Hello,
I need help solving this problem please.
In 1990 the expectancy of males in a certain country was 61.6 yrs. In 1996 it was 65.2 yrs let E represent the life expectancy in year t and let t represent the number of years since 1990.
E(t)=?t+?
Use function in 2003 E(13)=?

Answer by stanbon(75887)

In 1990 the expectancy of males in a certain country was 61.6 yrs. In 1996 it was 65.2 yrs
---
You have 2 points relating year and expectancy:
(0,61.6) and (6,65.2)
--------------------
let E represent the life expectancy in year t and let t represent the number of years since 1990.
---
slope = (65.2-61.6)/(6-0) = 3.6/6 = 0.6
intercept = E(0) = 61.6
------
E(t)=0.6t+61.6
-------
Use function in 2003
E(13)= 0.6*13+61.6
E(13) = 69.4
===============
Cheers,
Stan H.